OpenStudy (anonymous):

You are using 0.05 gallons per mile of an SUV gas tank that holds 15 gallons.

OpenStudy (anonymous):

what do u wanna no

OpenStudy (anonymous):



OpenStudy (anonymous):

Dimensional analysis; don't even need a calculator.\[.05\frac{gal}{mi}=\frac{1}{20}\frac{gal}{mi} \implies 20\frac{mi}{gal} \implies 20*15=300 mile range\]

OpenStudy (anonymous):

I think that's probably the question....